OpenStudy (anonymous):

Phillip bought 12 used CDs and DVDs. CDs cost $2 each and DVDs cost $3 each. He spent $31 not including tax. How many DVDs did Phillip buy?

OpenStudy (bahrom7893):

2*CD + 3*DVD = 31

OpenStudy (bahrom7893):

CD + DVD = 12

OpenStudy (bahrom7893):

Solve for CDs and DVDs

OpenStudy (anonymous):

Answers are: 10 7 5 2

OpenStudy (anonymous):

So... 2CD + 3DVD = 31 -2(CD + DVD = 12) -2CD + (-2DVD) = -24 DVD = 7 ?

OpenStudy (jhannybean):

why are you multiplying by -2?

OpenStudy (jhannybean):

You have two equations: x = CDs, y =DVDs. 1. 2x +3y = 31 2. x + y = 12 Take the second equation, x + y = 12, solve for 1, either x (CDs) or y (DVDs), and plug it into your first equation and solve for the first variable, then use that number to solve for the second variable.

OpenStudy (whpalmer4):

@jhannybean he confused you a bit with the notation — he multiplied the second equation by -2, thus eliminating the CD variable when they were combined. if done with your notation, \[2x+3y=31\]\[x+y=12\]multiply second eq. by -2 and combine \[2x+3y=31\]\[-2x-2y=-24\]---------------- \[y=7\]

OpenStudy (whpalmer4):

substitution is easy, but elimination can be even easier :-)

OpenStudy (jhannybean):

Truee...